A is a 2x2 matrix

Let A = {(1) (1/d) (c) (d)} Find the numbers c and d such that A^2 =0
1
Expert's answer
2021-08-04T18:13:42-0400
"A=\begin{pmatrix}
 1& 1/d \\
 c & d
\end{pmatrix}"

"A^2=\begin{pmatrix}
 1& 1/d \\
 c & d
\end{pmatrix}\begin{pmatrix}
 1& 1/d \\
 c & d
\end{pmatrix}"

"=\begin{pmatrix}
 1+c/d& 1+1/d \\
 c+cd & c/d+d^2
\end{pmatrix}=\begin{pmatrix}
 0 & 0 \\
 0 & 0
\end{pmatrix}"

"\begin{matrix}
1+c/d=0\\
1+1/d=0\\
c+c/d=0\\
c/d+d^2=0\\
\end{matrix}"

"c=1, d=-1"


"A=\begin{pmatrix}
 1& -1 \\
 1 & -1
\end{pmatrix}"


"A^2=\begin{pmatrix}
 1& -1 \\
 1 & -1
\end{pmatrix}\begin{pmatrix}
 1& -1 \\
 1 & -1
\end{pmatrix}=\begin{pmatrix}
 1-1& -1+1 \\
 1-1 & -1+1
\end{pmatrix}"

"=\begin{pmatrix}
 0& 0 \\
 0 & 0
\end{pmatrix}"